# A Hybrid Quantum-Classical Particle-in-Cell Method for Plasma Simulations

### v1

@@ 6 Conclusion and Discussion
In this work, we have proposed a hybrid quantum-classical PIC method for plasma simulations. In particular, we use a supervised quantum-classical hybrid neural network (HNN) to compute the electric potential given the charge densities, in other words, to solve the Poisson equation. We have tested our approach for the simulation of the two-stream instability test, a standard benchmark problem for the PIC codes. We have simulated the training data using linear solvers to compute the potentials for the charge densities at different time steps. We fed this data to train a classical neural network with three layers (CCC) and a HNN with classical-quantum-classical layers (CQC). The prediction accuracy, the phase-space evolution, and the energy conservation in the simulation suggest that a HNN (CQC) outperforms the classical model (CCC).
@@
We also investigated the usage of a physics-informed training approach for both the classical models and HNNs. Our tests indicate that the physics-informed approach can improve the learning strategies when it comes to predicting electric fields for an unseen charge density distribution. Here, we have mainly analyzed the errors in electric field predictions against the test solutions obtained from baseline PIC simulations. It is also important to investigate the error propagation in electric field prediction within the PIC loop and how the HNNs can be optimized to mitigate such propagation. Moreover, the proposed hybrid PIC method can also be generalized to solve Gauss’ law (.$\nabla.\mathrm{E}=\rho$) instead of the Poisson equation to obtain the electric field. In the classical electrostatic PIC methods, solving Gauss’ law is avoided as the results can be noisy, and it is not straightforward to define boundary conditions. However, in our proposed methods, such as data-driven or physics-informed approaches, we already provide the solutions at several grid points to train the neural network, which should be sufficient to obtain unique solutions of the equation. Further, Gauss’ law can be used to define a physics loss for the physics-informed approach.
@@
We have also analyzed the computational cost of training the neural networks. In particular, we have resorted to MPI-based data-parallelization scheme and explored the training performances with the varying number of parallel processes. A CQC model takes about 19x times that of the CCC model to complete the training, showing that long HNN training time on simulators is the major drawback of using quantum machine learning techniques at present. It could also be a possible direction to utilize accelerators such as GPUs to speed up quantum circuit simulations for the purpose of quantum machine learning, which could also allow us to study the performance scaling in the number of qubits and the variational ansatz depth. In addition, we plan to deploy our Quantum PIC on real quantum computer; however, we note that the algorithm requires a tight integration of quantum and computing parts, not available today quantum computing systems on the cloud.

### v2

@@ 6 Conclusion and Discussion
In this work, we have proposed a hybrid quantum-classical PIC method for plasma simulations. In particular, we use a supervised quantum-classical hybrid neural network (HNN) to compute the electric potential given the charge densities, in other words, to solve the Poisson equation. We have tested our approach for the simulation of the two-stream instability test, a standard benchmark problem for the PIC codes. We have demonstrated that HNNs can offer a performance advantage in solving Poisson equations—and, by extension, in simulating plasma behavior—with relatively fewer parameters compared to classical models, as shown through multiple experiments. However, it remains an open question how the inherent “quantumness” of the circuits systematically contributes to this performance gain [36]. A key challenge in applying QNNs to large-scale plasma simulations lies in understanding how performance scales with the dimensionality of the PDEs being solved. Additionally, it is important to investigate the trainability and expressivity of quantum circuits under realistic noise conditions, especially given the lack of efficient error correction protocols on current quantum hardware.
The other challenges of employing data-driven methods for simulating plasma include the availability of the data. The experimental techniques, such as Langmuir probes used to measure the electric potential within plasma, can be invasive and can have low spatial resolution [42]. We have shown that the usage of a physics-informed training approach can help overcome this disadvantage for both the classical models and HNNs. Our tests indicate that the physics-informed approach can improve the learning strategies when it comes to predicting electric fields for an unseen charge density distribution, and they are particularly effective in the presence of sparse data.
@@
Additionally, it is also important to investigate the error propagation in electric field prediction within the PIC loop and how the HNNs can be optimized to mitigate such propagation. Moreover, the proposed hybrid PIC method can also be generalized to solve Gauss’ law ($\nabla.\mathrm{E}=\rho$) instead of the Poisson equation to obtain the electric field. In the classical electrostatic PIC methods, solving Gauss’ law is avoided as the results can be noisy, and it is not straightforward to define boundary conditions. However, in our proposed methods, such as data-driven or physics-informed approaches, we already provide the solutions at several grid points to train the neural network, which should be sufficient to obtain unique solutions of the equation. Further, Gauss’ law can be used to define a physics loss for the physics-informed approach.
@@
We have also analyzed the computational cost of training the neural networks. In particular, we have resorted to MPI-based data-parallelization scheme and explored the training performances with the varying number of parallel processes. A CQC model takes about 19x times that of the CCC model to complete the training, showing that long HNN training time on simulators is the major challenge of testing and benchmarking quantum machine learning techniques at present. It could also be a possible direction to utilize accelerators such as GPUs to speed up quantum circuit simulations for the purpose of quantum machine learning, which could also allow us to study the performance scaling in the number of qubits and the variational ansatz depth [43].
Finally, we note that classical-quantum hybrid algorithms involve rapid exchange of large data between classical and quantum processors. This requires tightly-integrated classical and quantum computing parts, which allow low-latency workflows [44], a capability that is currently lacking in quantum systems accessible via the cloud. Achieving tight integration is a challenging task from both hardware and software points of view [45]. Advances in programming models and software tools aimed at overcoming these challenges [46, 43, 45] enhance the feasibility of deploying HNNs on heterogeneous classical–quantum systems. Such capabilities can, in turn, enable their use in computation-intensive tasks such as plasma PIC simulations, which is the topic of interest in this article.